A partition of a positive integer n is any way of writing n as a sum of one or more positive integers, in which we don't care about the order of the numbers in the sum. For example, the number 4 can be written as a sum of one or more positive integers ( which we don't care about the order of the numbers in the sum ) in exactly five ways:
4, 3+1, 2+2, 2+1+1, 1+1+1+1
So 4 has five partitions.
What is the number of partitions of the number 7?

Answer by jim_thompson5910(35256)

Adding up to 7

n = number of addends
Eg: n = 3 means we are adding 3 numbers to get to 7. The table shows 4 ways to add up 3 numbers to get to 7.


nNumber of unique waysSums
n=117
n=236+1,5+2,4+3
n=345+1+1,4+2+1,3+3+1,3+2+2
n=434+1+1+1,3+2+1+1,2+2+2+1
n=523+1+1+1+1,2+2+1+1+1
n=612+1+1+1+1+1
n=711+1+1+1+1+1+1

Note: the sums in red aren't special. They are done in red to provide alternating colors (red and black) so the sums stand out a bit better.